### v1

@@ Abstract
We introduce 1P1Q, a novel quantum data encoding scheme for high-energy physics (HEP), where each particle is assigned to an individual qubit, enabling direct representation of collision events without classical compression. We demonstrate the effectiveness of 1P1Q in quantum machine learning (QML) through two applications: a Quantum Autoencoder (QAE) for unsupervised anomaly detection and a Variational Quantum Circuit (VQC) for supervised classification of top quark jets. Our results show that the QAE successfully distinguishes signal jets from background QCD jets, achieving superior performance compared to a classical autoencoder while utilizing significantly fewer trainable parameters. Similarly, the VQC achieves competitive classification performance, approaching state-of-the-art classical models despite its minimal computational complexity. Furthermore, we validate the QAE on real experimental data from the CMS detector, establishing the robustness of quantum algorithms in practical HEP applications. These results demonstrate that 1P1Q provides an effective and scalable quantum encoding strategy, offering new opportunities for applying quantum computing algorithms in collider data analysis.
@@ I Introduction
The unprecedented collision energies achieved at latest- and next-generation colliders, like the Large Hadron Collider (LHC) and the Future Circular Collider (FCC), produce vast amounts of particle-level data, challenging the limits of conventional data analysis techniques in high energy physics (HEP). With the increasing complexity of these datasets, novel approaches that leverage cutting-edge computational paradigms have become indispensable. Quantum computing, with its ability to exploit superposition, entanglement, and interference, offers a promising framework to address some of the most intricate challenges in HEP.
@@
In recent years, Machine Learning (ML) has established itself as an indispensable tool for analysing HEP data, leading to significant advancements in tasks such as event classification, anomaly detection, and parameter estimation. The next frontier in this field is to extend these techniques to the quantum domain, leveraging Quantum Machine Learning (QML) to enhance data analysis capabilities. Any quantum machine learning algorithm consists of three key components: (1) the data encoding, which maps classical data onto quantum states; (2) the quantum model, typically implemented through quantum circuits and quantum operations such as entanglement between qubits; and (3) the loss function, whose optimization is crucial for training the quantum model. While substantial efforts have been devoted to designing quantum models and optimizing loss functions [1, 2], the choice of data encoding remains a critical yet underexplored aspect of QML for HEP applications. In this work, we propose a new encoding scheme tailored to HEP data, which we call 1P1Q (1 Particle - 1 Qubit). This encoding strategy assigns a separate qubit to each particle, enabling an efficient and direct representation of collider events without prior data compression.
@@
We demonstrate the effectiveness of quantum machine learning models in processing HEP data by employing two distinct approaches: a Quantum Autoencoder (QAE) [3] for unsupervised, unlabelled learning and a Variational Quantum Circuit (VQC) [4] for classification tasks. Both methods fully exploit the kinematic information of particles encoded via the 1P1Q scheme, using the quantum state representation to retain and process intricate correlations that are often lost in classical compression techniques. The QAE compresses quantum states by learning a lower-dimensional latent representation, making it well-suited for anomaly detection in collider physics, while the VQC employs a parametrized quantum circuit to discriminate between different class categories based on learned quantum features.
@@
We apply these quantum machine learning models to one of the most well-established use cases of classical machine learning in HEP: the discrimination of hadronically decaying resonances from the overwhelming background of QCD jets. Identifying such resonances, including those from top quarks, Higgs bosons, or hypothetical new physics states, is crucial for advancing our understanding of fundamental interactions. We find that for this task, VQC and QAE models acting on 1P1Q-encoded particle information result in highly performant QML algorithms, equal to or even better than comparable classical counterparts in the case of the QAE.
Compared to existing approaches [5, 6, 7, 8, 9], which first encode the information of the collider events into an abstract, typically compressed, latent representation using classical machine learning algorithms or domain-inspired high-level features, 1P1Q provides more immediate access to the raw information of collider events and allows for the direct exploitation of their content using quantum variational circuits. This has two main benefits: First, classical neural network architectures may result in a lossy representation and not fully capture the raw event content. Second, and related, 1P1Q provides a natural way of extending the input space as quantum computers and their simulators can accommodate more and more qubits.
@@
For the first time, our study explores QML in HEP using actual experimental data recorded by the CMS detector in 2016. This real-world application provides a critical test of the robustness and feasibility of QML strategies beyond simulated datasets. By demonstrating that QML models can be trained on and extract meaningful physics from real collision events, we show that 1P1Q can become a promising, lossless collider data encoding framework as quantum computing hardware advances.
@@
In this paper, we present the formulation of the 1P1Q encoding, explore its theoretical underpinnings, and discuss its potential applications to jet physics at the LHC and future colliders with the examples of anomaly detection and a supervised classifier. We demonstrate how this encoding captures the essential kinematic features of jets and showcase its utility in leveraging quantum algorithms to analyze jet substructure, and compare to state-of-the-art classical machine learning algorithms.
II 1P1Q - Particle Encoding
@@
The 1P1Q method directly encodes the kinematics of a particle on a qubit. The pseudorapidity $\eta$ and the azimuthal angle $\phi$ of the particle, modulated by the transverse momentum $p_{\mathrm{T}}$ normalized to the $p_{\mathrm{T}}$ of the jet, are used as spherical coordinates on the Bloch sphere to orient the qubit, enabling a compact and information-rich representation of each particle as a quantum state that is not dependent on the energy scale of the jet, thus facilitating a straightforward transition to the quantum domain. These feature encodings, represented by rotation angles about the $Y$ and $X$ axes respectively, are then additionally scaled by a factor $f=f(w)$ constrained to lie between $[1,2\pi+1]$, where $w$ is a trainable parameter. This ensures the particles can spread out across the Bloch sphere instead of clustering too closely around its North Pole. This is especially true when using normalized $p_{\mathrm{T}}$ distributions, where each particle $p_{\mathrm{T}}$ gets normalized by the jet $p_{\mathrm{T}}$ and thus tends to have small values, as can be seen in Figure 1, where we show the effect of this scaling for the ten hardest particles from a $QCD$ jet and a jet produced from a top quark decay. The encoding can then be summarized as:
$f\cdot\dfrac{p_{\mathrm{T}}}{p_{\mathrm{T,jet}}}\cdot(\eta-\eta_{jet})\quad$ $\to\quad\theta$ (1)
$f\cdot\dfrac{p_{\mathrm{T}}}{p_{\mathrm{T,jet}}}\cdot(\phi-\phi_{jet})\quad$ $\to\quad\varphi$ (2)
$(p_{\mathrm{T}},\eta,\phi)\quad$ |$\to\quad\ket{\psi}=R_{X}(\varphi)R_{Y}(\theta)\ket{0}$
$~{}~{}~{}~{}=\alpha(\theta,\varphi)\ket{0}+\beta(\theta,\varphi)\ket{0}$ (3)
@@
In the above equations, we take the coordinates $\eta$ and $\phi$ relative to the jet axis. While LHC collisions can result in hundreds of final-state particles, we limit our studies to the intrinsic structure of jets rather than entire events due to the limitations of currently available quantum computers and their simulators. Studying the substructure of jets can be a very powerful way to disentangle boosted, hadronically decaying electroweak-scale or beyond-the-standard-model resonances from QCD-induced jets [10]. Furthermore, in order to allow for efficient simulation of our quantum circuits on a simulator, we limit our study by using up to ten hardest jet constituents, which are expected to carry most of the information relevant to jet substructure analysis. Fig. 1 shows an example of the encoding for the ten hardest particles in a top-decay event (red) and a QCD light-flavour jet event (blue). Each particle is encoded on a different qubit. The scaling factor in Eq. 4 increases the angles on the Bloch sphere.
Figure 1: Bloch Sphere: Effect of the input scaling described in Equation 4 when applied to the ten hardest particles of a QCD and top jet with comparable jet $p_{\mathrm{T}}$. The value of $w$ converges to a scaling factor $f=7.268$ for the VQC.
III Quantum circuits for anomaly detection and classification
Anomaly detection is pivotal in high-energy physics, where new physics signatures often manifest as deviations from Standard Model predictions. The vast data volumes and model-agnostic nature of potential signals require robust detection methods. In classical machine learning, unsupervised learning models such as autoencoders have emerged as efficient tools for anomaly detection [11]. These work by compressing the input data into a latent space and reconstructing it, highlighting anomalous events through elevated reconstruction errors.
Quantum autoencoders (QAEs) extend this paradigm to quantum devices [3]. A QAE consists of an encoder and a decoder implemented via variational quantum circuits. The encoder compresses the input quantum state into a smaller latent representation, discarding qubits as necessary. The decoder then reconstructs the input from this latent state using the Hermitian conjugate of the encoding operators.Without sufficient dimensionality reduction in the latent space, a network would be able to learn the identity transformation trivially. Unlike classical autoencoders, QAEs perform this reduction by replacing the discarded qubits (hereafter referred to as trash states) with reference states that are initialized to |$\ket{0}$, thus creating an information bottleneck. Naturally, this would require the number of trash states $N_{\text{trash}}$ and reference states $N_{\text{ref}}$ to be equal.
@@ III Quantum circuits for anomaly detection and classification
The unitary transformations within a QAE consist of a combination of parameterized single-qubit rotations and multi-qubit entangling gates, such as the Controlled NOT (CNOT) gate. To demonstrate its applicability to the task of anomaly detection, we use the procedure described in Equations 1 and 2, to encode information into our quantum circuit. Next, to allow the network to learn higher-order non-linear terms, we entangle all possible pairs of qubits using two-qubit CNOT gates. For a system of $N$ qubits, this would require a total of $N(N-1)/2$ operations. Finally, to ensure the network learns an optimal representation of the input space, we apply three parameterized rotations, one along each axis to each qubit. These trainable parameters are optimized using the classical Adam Optimizer [12], with a scheduler that periodically decays the learning rate. The entanglement and rotation operations are summarized in Equation 5.
$U(\Theta)=\left(\bigotimes_{i=1}^{N}R_{X}(\phi_{i})R_{Z}(\theta_{i})R_{Y}(\omega_{i})\right)\otimes\left(\bigotimes_{1\leq i<j\leq N}C_{ij}\right).$ (5)
The quantum circuits presented in this study are simulated and optimized using the Quantum Machine Learning library pennylane[[13]] with the lightning.gpu and lightning.kokkos devices. To train the QAE, we seek to maximize the fidelity between the output and input states in the subspace relevant for reconstruction. Using an equivalent approach first introduced in [14], we minimize instead the cost function defined as the negative of the fidelity between the trash and reference states. This fidelity measurement is performed using a SWAP test [15], which requires an ancillary qubit initialized to |$\ket{0}$.
Using specific examples, QAEs have shown first successes compared to classical AEs regarding training efficiency and performance. The QAE of [3] requires fewer training samples to converge and generally outperforms CAEs in anomaly detection tasks.
In addition to its applicability for as an anomaly detection tool, we demonstrate that the 1P1Q encoding can just as well be used for classification within a supervised learning paradigm. For this task, we use a Variational Quantum Classifier (VQC) that learns to separate signal and background classes [4], in this case jets initiated by top quarks decaying hadronically versus jets initiated by light quarks or gluons.
The VQC is similar to the encoder of the QAE, with features being first encoded into the circuit using the procedure outlined in Equations 1-4. This is followed by entanglement operations using two-qubit CNOT gates between adjacent qubits, to allow for the introduction of non-linear terms into the network. For an input space of $N$ qubits, this would require $N+1$ CNOT operations. Lastly, three trainable rotation gates are applied to each qubit, which allows the network to learn an optimal representation that separates signal from background, as summarized in Equation 6. Unlike the autoencoder, which is optimized by maximizing the fidelity between the trash and reference states, the VQC is optimized by performing a measurement of the expectation value of the Pauli $Z$ observable on a target qubit, usually the first qubit, which is bounded in $[-1,1]$. The final prediction from the VQC (Equation 7) is arrived at by adding a classical bias term, also a trainable parameter, to this expectation value. The circuit optimization is performed using the Mean Squared Error (MSE) between the prediction and the truth label as the loss function, along with the Adam Optimizer. The complete VQC circuit is shown in Figure 2 for an input of $N=8$ particles.
$U(\Theta)=$ $\left(\bigotimes_{i=1}^{N-1}R_{X}(\phi_{i})R_{Z}(\theta_{i})R_{Y}(\omega_{i})\right)$
$\otimes\left(\bigotimes_{i=0}^{N}C_{i,(i+1)\bmod N}\right)$ (6)
$f(x)$ $=\langle q(x)|Z|q(x)\rangle+b$ (7)
We highlight that for an input space comprising of the $N$ hardest input particles per jet, the QAE requires only $\mathbf{3N+1}$ trainable parameters to learn a suitable reconstruction of the inputs. Likewise, the VQC requires only $\mathbf{3N+2}$ parameters.
IV Training Datasets
@@
The JetClass dataset introduced in [16], is a vast library containing $100$M jets for training, $5$M jets for validation and $20$M jets for inference respectively, equally divided into ten classes. For the purpose of training our QAE, we use jets initiated by light quarks or gluons, hereafter referred to as the background jets. Such jets are produced in abundance at the LHC. Inference is then performed on both background as well as signal jets, comprising jets produced by decays of particles such as the Higgs, W or Z bosons, or the top quark. Since the QAE is trained to reconstruct only jets initiated by light quarks or gluons, we expect this reconstruction quality, defined as the fidelity between the trash and reference states described in Section III, to differ between signal and background, thus providing a degree of separation.
$F$ $=\sum_{i=1}^{N_{\text{ref}}}\braket{T_{i}}{R_{i}}$ $\forall~{}~{}~{}~{}T_{i}\in H^{T},R_{i}\in H^{\text{ref}}$ (8)
For training the QAE and VQC respectively, we first sample jets such that each class has a flat distribution in jet $p_{\mathrm{T}}$, in the range $[500,1000]$ GeV, so as to not bias the training towards the scale of the jet, allowing us to purely focus on jet substructure.
We also demonstrate the effectiveness of our encoding for anomaly detection by training the QAE on real world data, recorded with the CMS Detector at the LHC [17] in 2016. This dataset, recently released in a machine-learning friendly format [18] is dominated by QCD jets, with contamination from other sources being estimated at less than $1\%$ overall.
V Results and Benchmarking
V.1 Quantum Autoencoder
The QAE is an example of an unsupervised learning algorithm, since it does not require labelled data and can be trained only on the background class, which in this case comprises jets initiated by light quarks or gluons. We use a total of $10000$ events for training and $2500$ events for validation. Inference is performed on $10000$ events of each of the following class of events: $H\to b\overline{b}$, $H\to c\overline{c}$, $H\to gg$, $W\to q\overline{q}$, $Z\to q\overline{q}$ and $t\to bq\overline{q\prime}$. The signal jets are also sampled to have a flat $p_{\mathrm{T}}$ distribution in the range $[500,1000]$ GeV.
@@ V.1 Quantum Autoencoder
Figure 3: Quantum Fidelity distributions for a system with $10$ input qubits and a latent space of $2$ qubits
@@
We train the QAE on two different sets: simulated QCD jets from the JetClass dataset, and the real jet data recorded with the CMS Detector in 2016. Figure 3 shows that a QAE trained on data is comparable in performance to one trained on simulated events, when applied to the task of differentiating between simulated QCD events and simulated signal models.
For the first time, to our knowledge, we also show the results in Figure 3 of using a quantum machine learning algorithm with CMS open data to prove that the QAE can learn the underlying physics of the jet and is not biased by the detector response.
@@
Figure 4 illustrates the AUC scores as a function of the quantum fidelity metric $\langle 1-\text{Fidelity}\rangle_{\text{QCD}}$ for different quantum autoencoder (QAE) configurations trained on either simulated JetClass QCD jets (dashed lines) or real CMS data (dotted lines). The three panels correspond to different input sizes, namely 6, 8, and 10 jet constituents, while different marker shapes indicate varying latent space sizes. Across all configurations, higher AUC scores correlate with increased $\langle 1-\text{Fidelity}\rangle_{\text{QCD}}$ , indicating that the fidelity loss serves as a useful proxy for anomaly detection performance. The performance is particularly strong for top quark jets ($t\to bqq{\prime}$, red), which consistently achieve the highest AUC scores, followed by Higgs decays to bottom quarks ($H\to b\overline{b}$, blue). The similarity in trends between models trained on real and simulated data highlights the robustness of the QAE in learning fundamental jet substructure features independent of dataset origin. Additionally, larger input sizes tend to improve performance, suggesting that incorporating more jet constituents provides richer representations for classification.
To benchmark the performance of the 1P1Q-encoded QAE against a classical counterpart, we trained on simulated QCD jets and considered as anomalies to this background the signals of hadronic $W$ boson, Higgs boson and top quark decays. While the QAE has a simple structure of 10 input qubits, followed by the circuit of Figure 2 (left) and a latent space of 2, the classical autoencoder (CAE) model is allowed to be significantly larger, consisting of an encoder model containing an input feature vector of size 30, to be able to encode the same number of features encoded on the QAE, followed by dense layers of size $20-16-12$ and a latent space 6. The decoder of the CAE has an identical, yet inverse, structure.
The QAE achieves superior performance compared to the CAE across all signal types. The QAE maintains this advantage despite having only 32 trainable parameters, in contrast to the CAE’s 2,500 parameters, demonstrating the potential of quantum machine learning to capture relevant physics with a significantly reduced model complexity efficiently.
Algorithm Signals $W\to q\overline{q}$ $H\to b\overline{b}$ $t\to bq\overline{q^{\prime}}$
QAE 0.693 0.757 0.861
@@
Table 1: AUC scores QAE vs CAE: trained on a 10-particle input space
V.2 Variational Quantum Circuit
For the supervised classification task, we employ the VQC to distinguish jets originating from top quark decays ($t\to bq\overline{q}^{\prime}$) from those initiated by light quarks or gluons. The VQC is trained using 1,000 events and validated on 500 events, with an inference dataset of 10,000 events equally distributed between signal and background classes, with all jets sampled to have a flat $p_{\mathrm{T}}$distribution.
Figure 4 presents the Receiver Operating Characteristic (ROC) curves comparing the VQC trained on the 1P1Q-encoded dataset against the state-of-the-art Particle Transformer (ParT) classifier[19], both trained using the 10 hardest particles per jet, and the same number of jets. The ROC curve demonstrates that the VQC achieves strong classification performance, with an AUC score of 0.876. Although the Particle Transformer model achieves a slightly higher AUC of 0.898, the VQC remains competitive despite having significantly fewer trainable parameters (32 vs. over 2 million). Furthermore, at a signal efficiency of 99%, the VQC achieves a background rejection of REJ99 = 2.288, outperforming the Particle Transformer’s REJ99 = 2.048.
@@ V.2 Variational Quantum Circuit
V.3 1P1Q encoding
As detailed in Section I, quantum machine learning methods consist not only of the models, such as QAE or VQC, but also of the encoding step that maps classical data onto a quantum device. Both components jointly define the expressivity and efficiency of a quantum machine learning approach. Thus, to assess the QAE and VQC models’ ability to utilize the information provided by the 1P1Q data encoding, we successively reduce the features encoded on each qubit and observe the resulting performance degradation.
@@
VQC 0.876 0.846 0.805 0.806
QAE 0.861 0.813 0.791 0.804
Table 2: AUC scores vs Input Features for the VQC and QAE, for the benchmark signal: $t\to bq\overline{q\prime}$
@@
VI Summary
In this work, we introduced the 1P1Q encoding scheme, a novel approach for representing particle physics data on quantum hardware by assigning each particle to a separate qubit. This encoding allows direct utilization of raw collision event data without classical compression, enhancing the potential expressivity and efficiency of QML models. We demonstrated the effectiveness of this approach using two quantum machine learning models: a QAE for anomaly detection and a VQC for supervised classification. These models were trained on both simulated and real experimental data, establishing the robustness of quantum algorithms in high-energy physics applications.

### v2

@@ Abstract
We introduce 1P1Q, a novel quantum data encoding scheme for high-energy physics (HEP), where each particle is assigned to an individual qubit, enabling direct representation of collision events on quantum circuits without classical compression. We demonstrate the effectiveness of 1P1Q in quantum machine learning (QML) through two applications: a Quantum Autoencoder (QAE) for unsupervised anomaly detection and a Variational Quantum Circuit (VQC) for supervised classification of top quark jets. Our results show that the QAE successfully distinguishes signal jets from background QCD jets, achieving superior performance compared to a classical autoencoder while utilizing significantly fewer trainable parameters. Similarly, the VQC achieves competitive classification performance, approaching state-of-the-art classical models despite its minimal computational complexity. Furthermore, we validate the QAE on real experimental data from the CMS detector, establishing the robustness of quantum algorithms in practical HEP applications. These results demonstrate that 1P1Q provides an effective and scalable quantum encoding strategy, offering new opportunities for applying quantum computing algorithms in collider data analysis.
@@ Introduction.
The unprecedented collision energies achieved at latest- and next-generation colliders, like the Large Hadron Collider (LHC) and the Future Circular Collider (FCC), produce vast amounts of particle-level data, challenging the limits of conventional data analysis techniques in high energy physics (HEP). With the increasing complexity of these datasets, novel approaches that leverage cutting-edge computational paradigms have become indispensable. Machine Learning (ML) has established itself as an indispensable tool for analysing HEP data in recent years, leading to significant advancements in tasks such as event classification, anomaly detection, and parameter estimation. An emerging frontier in this field is to extend these techniques to the quantum domain, leveraging Quantum Machine Learning (QML) to enhance data analysis capabilities. Quantum computing, with its ability to exploit superposition, entanglement, and interference, offers a promising framework to address some of the most intricate challenges in HEP [1, 2, 3].
@@
Any quantum machine learning algorithm consists of three key components: (1) the data encoding, which maps classical data onto quantum states; (2) the quantum model, typically implemented through quantum circuits and quantum operations such as entanglement between qubits; and (3) the loss function, whose optimization is crucial for training the quantum model. While substantial efforts have been devoted to designing quantum models and optimizing loss functions [4, 5, 6, 7, 8, 9, 10, 11, 12, 13], the choice of data encoding remains a critical yet underexplored aspect of QML for HEP applications. In this work, we propose a new encoding scheme tailored to HEP data, which we call 1P1Q (1 Particle - 1 Qubit). This encoding strategy assigns a separate qubit to each particle, enabling an effective representation of collider events within quantum circuits without prior data compression. Establishing a direct correspondence between particles and qubits is warranted to advance the usefulness of quantum computing for HEP in particular upon entering an era in which crucial LHC measurements reveal the quantum mechanical nature of the imprints left behind by the collisions, such as the recently observed entanglement in top quark pair production [14, 15].
@@
We demonstrate the effectiveness of 1P1Q-based QML models in processing HEP data by employing two distinct approaches: a Quantum Autoencoder (QAE) [16] for unsupervised, unlabelled learning and a Variational Quantum Circuit (VQC) [17] for classification tasks. The QAE compresses quantum states by learning a lower-dimensional latent representation, making it well-suited for anomaly detection in collider physics, while the VQC employs a parametrized quantum circuit to discriminate between different class categories based on learned quantum features. Both methods fully exploit the kinematic information of particles encoded via the 1P1Q scheme, using the quantum state representation to retain and process intricate correlations between particles. Such information is at risk of being lost in existing approaches [18, 19, 20, 21, 22, 23, 24, 25], which first encode the information of the collider events into an abstract, typically compressed, latent representation using classical machine learning algorithms or domain-inspired high-level features. Instead, 1P1Q guarantees more immediate access to the raw information of collider events and allows for the direct exploitation of their content using quantum variational circuits. 1P1Q therefore also provides a natural way of extending the input space as quantum computers and their simulators can accommodate more and more qubits.
@@
We apply the QML models to one of the most well-established use cases of classical machine learning in HEP: the discrimination of hadronically decaying resonances from the overwhelming background of QCD jets. Identifying such resonances, including those from top quarks, Higgs bosons, or hypothetical new physics states, is crucial for advancing our understanding of fundamental interactions. We find that for this task, VQC and QAE models acting on 1P1Q-encoded particle information result in highly performant QML algorithms, equal to or even better than comparable classical counterparts in the case of the QAE.
@@
For the first time, our study explores QML in HEP using actual experimental data recorded by the CMS detector [26] in 2016. This real-world application provides a critical test of the robustness and feasibility of QML strategies in general, and 1P1Q in particular, beyond simulated datasets. By demonstrating that QML models can be trained on and extract meaningful physics from real collision events, we show that 1P1Q can become a modern, lossless collider data encoding framework as quantum computing hardware advances.
@@
In the following, we present the formulation of the 1P1Q encoding, explore its theoretical underpinnings, and discuss its potential applications to jet physics at the LHC and future colliders with the examples of anomaly detection and a supervised classifier. We demonstrate how this encoding captures the essential kinematic features of jets and showcase its utility in leveraging quantum algorithms to analyze jet substructure, and compare to state-of-the-art classical machine learning algorithms.
1P1Q - Particle Encoding.
@@
The 1P1Q method directly encodes the kinematics of a particle on a qubit. The pseudorapidity $\eta$ and the azimuthal angle $\phi$ of the particle, modulated by the transverse momentum $p_{\mathrm{T}}$ normalized to the $p_{\mathrm{T}}$ of the jet, are used as spherical coordinates on the Bloch sphere to orient the qubit, enabling a compact and information-rich representation of each particle as a quantum state that is not dependent on the energy scale of the jet, thus facilitating a straightforward transition to the quantum domain. These feature encodings, represented by rotation angles about the $Y$ and $X$ axes respectively, are then additionally scaled by a factor $f=f(w)$ constrained to lie between $1,2\pi+1]$, where $w$ is a trainable parameter. This ensures the particles can spread out across the Bloch sphere instead of clustering too closely around its North Pole. The encoding can then be summarized as:
$f\cdot\dfrac{p_{\mathrm{T}}}{p_{\mathrm{T,jet}}}\cdot(\eta-\eta_{\text{jet}})\quad$ $\to\quad\theta$ (1)
$f\cdot\dfrac{p_{\mathrm{T}}}{p_{\mathrm{T,jet}}}\cdot(\phi-\phi_{\text{jet}})\quad$ $\to\quad\varphi$ (2)
[…]
@@
In the above equations, we take the coordinates $\eta$ and $\phi$ relative to the jet axis. While LHC collisions can result in hundreds of final-state particles, we limit our studies to the intrinsic structure of jets rather than entire events due to the limitations of currently available quantum computers and their simulators. Studying the substructure of jets can be a very powerful way to disentangle boosted, hadronically decaying electroweak-scale or beyond-the-standard-model resonances from QCD-induced jets [27]. Furthermore, in order to allow for efficient simulation of our quantum circuits on a simulator, we limit our study by using up to ten hardest jet constituents, which are expected to carry most of the information relevant to jet substructure analysis. Figure 1 shows an example of the encoding for the ten hardest particles in a jet from a top quark decay (red) and in a QCD light-flavor jet (blue). Each particle is encoded on a different qubit. As can be seen, the encoding is rooted in the kinematic structure of jets, mapping angular coordinates and transverse momentum to the Bloch sphere to preserve geometric and energy correlations among constituents. The scaling function $f$ in Eq. 4 increases the angles on the Bloch sphere.
Figure 1: Bloch Sphere: Effect of the input scaling described in Eq. 4 when applied to the ten hardest particles of a QCD and top jet with comparable jet $p_{\mathrm{T}}$. The value of $w$ converges to a scaling factor $f=7.268$ for the VQC.
Anomaly detection and classification with 1P1Q.
Anomaly detection has emerged as a pivotal tool in model-agnostic searches for new physics, where the only assumption is that deviations from Standard Model signatures are present [28]. Using classical machine learning, unsupervised learning models based on autoencoders are employed as efficient tools for anomaly detection [29, 30, 31]. These work by compressing the input data into a lower-dimensional latent space and attempting to reconstruct the input, highlighting anomalous events through elevated reconstruction errors.
[…]
@@
To demonstrate the applicability of 1P1Q in a QAE, we use Eqs. 1 and 2 to encode information into a QAE with an architecture following the left circuit of Fig. 2. For illustration purposes, the figure shows a $4$-particle-input example, while the QAEs trained are using $6$, 88, and $10$ particles as input, respectively. To allow the network to learn higher-order non-linear terms, we entangle all possible pairs of qubits using two-qubit CNOT gates. To ensure the network learns an optimal representation of the input space, we apply three parameterized rotations, one along each axis of each qubit. The entanglement and rotation operations are summarized in Eq. 5.
$U(\Theta)=\left(\bigotimes_{i=1}^{N}R_{X}(\phi_{i})R_{Y}(\theta_{i})R_{Z}(\omega_{i})\right)\otimes\left(\bigotimes_{1\leq i<j\leq N}C_{ij}\right).$ (5)
@@
The QAE is trained to reconstruct jets initiated by light quarks or gluons. We expect this reconstruction quality, defined as the fidelity between the trash and reference states, to differ between signal and background, thus providing a degree of separation.
In addition, we demonstrate that the 1P1Q encoding can just as well be used for classification within a supervised learning paradigm. For this task, we use a VQC [17] that learns to separate signal and background classes, in this case jets from hadronically decaying top quarks versus jets initiated by light quarks or gluons.
The VQC architecture is similar to that of the QAE, with features being first encoded into the circuit using the procedure outlined in Eqs. 1-4, followed by entanglement operations using two-qubit CNOT gates to allow for the introduction of non-linear terms into the network. Only adjacent qubits are entangled for the VQC, which we find gives the best performance. Three trainable rotation gates are applied to each qubit, as summarized in Eq. 6. Unlike the QAE, which is optimized by maximizing the fidelity between the trash and reference states, the VQC is optimized by performing a measurement on the first qubit which is bound in [1,1][-1,1] and comparing it to the ground truth label. The VQC circuit is shown in Fig. 2 for an input of $N=8$ particles.
$U(\Theta)=$ $\left(\bigotimes_{i=1}^{N}R_{X}(\phi_{i})R_{Y}(\theta_{i})R_{Z}(\omega_{i})\right)$
$\otimes\left(\bigotimes_{i=1}^{N}C_{i,(i+1)\bmod N}\right)$ (6)
We highlight that for an input space comprising of the $N$ hardest input particles per jet, the QAE requires only $\mathbf{3N+1}$ trainable parameters to learn a suitable reconstruction of the inputs. Likewise, the VQC requires only $\mathbf{3N+2}$ parameters (the extra parameter of the VQC being a trainable bias term that is added to the measurement).
Results and Benchmarking.
@@
Figure 3: Quantum fidelity distributions for a system with 1010 input qubits and a latent space of 22 qubits.
@@
We train the QAE on two different sets: simulated QCD jets from the JetClass dataset [32], and real jet data recorded with the CMS Detector in 2016 [33], which is >>99% pure in QCD jets. For the first time, to our knowledge, Fig. 3 shows that a QAE trained on real experimental data is comparable in performance to one trained on simulated events, when applied to the task of differentiating between simulated QCD events and simulated signal models. This suggests that a QAE using 1P1Q encoding, despite not using high-level information, can extract the underlying physics of the jet based on the individual particles encoded onto qubits and is neither corrupted nor biased by the detector response.
@@
Figure 4 illustrates the AUC scores as a function of the quantum fidelity metric $\langle 1-\text{Fidelity}\rangle_{\text{QCD}}$ for different QAE configurations trained on either simulated JetClass QCD jets (dashed lines) or real CMS data (dotted lines). The three panels correspond to different input sizes, namely 6, 8, and 10 jet constituents, respectively, while different marker shapes indicate varying latent space sizes. Across the various configurations, we observe a trend that higher AUC scores correlate with increased $\langle 1-\text{Fidelity}\rangle_{\text{QCD}}$ , indicating that the fidelity loss serves as a useful proxy for anomaly detection performance. The discrimination is particularly strong for top quark jets ($t\to bq\overline{q^{\prime}}$, red), which consistently achieve the highest AUC scores, followed by Higgs decays to bottom quarks ($H\to b\overline{b}$, blue). The similarity in trends between models trained on real and simulated data highlights the robustness of the QAE in learning fundamental jet substructure features independent of dataset origin. Additionally, larger input sizes tend to improve performance, suggesting that incorporating more jet constituents provides richer representations for classification.
To benchmark the performance of the 1P1Q-encoded QAE against a classical counterpart, we trained on simulated QCD jets and considered as anomalies to this background the signals of hadronic $W$ boson, Higgs boson and top quark decays. While the QAE has a simple structure of 10 input qubits, followed by the circuit of Fig. 2 (left) and a latent space of 2, the classical autoencoder (CAE) model is allowed to be significantly larger, consisting of an encoder model containing an input feature vector of size 30, to be able to encode the same number of features encoded on the QAE, followed by dense layers of size 20161220-16-12 and a latent space 6. The decoder of the CAE has an identical, yet inverse, structure.
The QAE achieves superior performance compared to the CAE across all signal types. The QAE maintains this advantage despite having only 31 trainable parameters, in contrast to the CAE’s $\sim$2,500 parameters, demonstrating the potential of quantum machine learning to capture relevant physics with a significantly reduced model complexity efficiently.
Table 1: AUC scores for QAE vs CAE, trained on a 10-particle input space.
Algorithm Signals $W\to q\overline{q^{\prime}}$ $H\to b\overline{b}$ $t\to bq\overline{q^{\prime}}$
QAE 0.715 0.774 0.872
@@
For the supervised classification task, we employ the VQC to distinguish jets originating from top quark decays ($t\to bq\overline{q^{\prime}}$) from those initiated by light quarks or gluons. Figure 5 presents the Receiver Operating Characteristic (ROC) curves comparing the VQC trained on the 1P1Q-encoded dataset against the state-of-the-art Particle Transformer (ParT) classifier [34], both trained using the 10 hardest particles per jet, and the same dataset of 1000 jets. The ROC curve demonstrates that the VQC achieves strong classification performance, with an AUC score of 0.885. Although the ParT model achieves a slightly higher AUC of 0.898, the VQC remains competitive despite having significantly fewer trainable parameters (32 vs. over 2 million). Furthermore, for high signal efficiencies (>>80%), the VQC yields better background rejection than the Particle Transformer.
@@
1P1Q encoding.
The expressivity and efficiency of a quantum machine learning approach are jointly defined both by the model (e.g., QAE or VQC) and the encoding step that maps classical data onto a quantum device. Thus, to assess the models’ ability to utilize the information provided by the 1P1Q data encoding, we successively reduce the features encoded on each qubit and observe the resulting performance degradation.
Table 2: AUC scores vs input features for the VQC and QAE, for the benchmark signal $t\to bq\overline{q^{\prime}}$.
@@
VQC 0.886 0.856 0.808 0.857
QAE 0.872 0.825 0.823 0.827
@@
Summary.
In this work, we introduced the 1P1Q encoding scheme, a novel approach for representing particle physics data on quantum hardware by assigning each particle to a separate qubit. This encoding allows direct utilization of raw collision event data without classical compression, enhancing the potential expressivity and efficiency of QML models. We demonstrated the effectiveness of this approach using two quantum machine learning models: a QAE for anomaly detection and a VQC for supervised classification.